Skip to main contentSkip to solution

If a fair coin is tossed 10 times, then the probability of obtaining at least one head is :

Solution

✅ Correct Option: 3

P(at least one head)=1−P(no head)=1−(1/2)10=1−1/1024=1023/1024P(\text{at least one head}) = 1 - P(\text{no head}) = 1 - (1/2)^{10} = 1 - 1/1024 = 1023/1024.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question